A mixture of 16% disinfectant solution is to be made from 20% and 14% disinfectant solutions. How much of each solution should be used in 15 gallons of the 16% solution is needed.
Answer by checkley77(12844)

.20x+.14(15-x)=.16*15
.20x+2.1-.14x=2.4
.06x=2.4-2.1
.06x=.3
x=.3/.06
x=5 gal. of 20% solution is used.
15-5=10 gal. or 14% solution is used.
Proof:
.20*5+.14*10=.16*15
1+1.4=2.4
2.4=2.4